# mnsymbols’102 mnsymbols’107 mnsymbols’164 mnsymbols’171

### v1

@@ Abstract.
In this paper, we propose TANQ-Sim, a full-scale density matrix based simulator designed to simulate practical deep circuits with both coherent and non-coherent noise. To address the significant computational cost associated with such simulations, we propose a new density-matrix simulation approach that enables TANQ-Sim to leverage the latest double-precision tensorcores (DPTCs) in NVIDIA Ampere and Hopper GPUs. To the best of our knowledge, this is the first application of double-precision tensorcores for non-AI/ML workloads. To optimize performance, we also propose specific gate fusion techniques for density matrix simulation. For scaling, we rely on the advanced GPU-side communication library NVSHMEM and propose effective optimization methods for enhancing communication efficiency. Evaluations on the NERSC Perlmutter supercomputer demonstrate the functionality, performance and scalability of the simulator. We also present three case studies to showcase the practical usage of TANQ-Sim, including teleportation, entanglement distillation, and Ising simulation. TANQ-Sim will be released on GitHub.
@@ 5.2. Benchmark Results
Performance Scaling: Figure 7 shows the performance scaling of simulation time for the circuit benchmarks on the Perlmutter HPC. We scale from 2 nodes, 8 GPUs to 512 nodes, 2,048 GPUs. For circuits with small numbers of qubits, such as adder, qpe and vqe, some bars are missing on the right, as to run on the DPTUs, we will need at least 2 qubits per GPU to perform a 2-qubit gate. On the other hand, for circuits with larger number of qubits, such as dnn and qft, an insufficient number of GPUs cannot accommodate the density matrix in the SHMEM space, which is 16GB per GPU. Remind that for density-matrix simulation, one more qubit costs 4x GPU memory cost.
From the figure, it can be seen that: (i) TANQ-Sim using NVSHMEM for inter-node communication can scale to 512 nodes, 2,048 GPUs without any issues; (ii) For circuits with less qubits, such as adder, qpe and vqe, adding more GPUs does not improve but essentially degrade performance, as GPUs are consistently underutilized with expanded communication and synchronization overhead. (iii) For circuits with more qubits to saturate each GPUs, the performance continuously improves with extra GPUs in scaling.
Figure 7. TANQ-Sim performance scaling on Perlmutter.

### v2

@@ Abstract.
In this paper, we propose TANQ-Sim, a full-scale density matrix based simulator designed to simulate practical deep circuits with both coherent and non-coherent noise. To address the significant computational cost associated with such simulations, we propose a new density-matrix simulation approach that enables TANQ-Sim to leverage the latest double-precision tensorcores (DPTCs) in NVIDIA Ampere and Hopper GPUs. To the best of our knowledge, this is the first application of double-precision tensorcores for non-AI/ML workloads. To optimize performance, we also propose specific gate fusion techniques for density matrix simulation. For scaling, we rely on the advanced GPU-side communication library NVSHMEM and propose effective optimization methods for enhancing communication efficiency. Evaluations on the NERSC Perlmutter supercomputer demonstrate the functionality, performance, and scalability of the simulator. We also present three case studies to showcase the practical usage of TANQ-Sim, including teleportation, entanglement distillation, and Ising simulation. TANQ-Sim will be released on GitHub.
@@ 5.2. Benchmark Results
Performance Scaling: As quantum circuit simulation overhead increases exponentially with the number of qubits being simulated, achieving strong scaling performance across larger numbers of compute nodes is crucial for handling larger circuit simulations.
Table 4. Larger Benchmark Circuits
Circuit Algorithm Qubits Gates Density Matrix Size
DNN Machine Learning 16 1041 128 GB
QFT Hidden Subgroup 16 611 128 GB
QV Quantum Volume 17 36 512 GB
QEC Surface Code 17 54 512 GB
Adder Quantum Arithmetic 18 277 2 TB
SQRT Quantum Arithmetic 18 2143 2 TB
RCS Random Circuit Sampling 19 137 8 TB
BV Hidden Subgroup 19 57 8 TB
QRAM Quantum Memory 20 308 32 TB
QAOA Search and Optimization 20 101 32 TB
GHZ Entanglement Generation 21 43 128 TB
W_State Entanglement Generation 21 103 128 TB
To evaluate this scaling, we tested larger benchmark circuits shown in Table 4 with 16 to 21 qubits. Notably, the largest benchmark circuits require 128 TB of memory. For these tests, we employed a larger device noise profile - a 30-qubit device with similar gate and decoherence error rates compared to the 16-qubit ibmq_guadalupe device used previously. This device features all-to-all qubit connectivity and it is chosen to prevent the introduction of additional ancilla qubits during circuit transpilation.
Figure 7. TANQ-Sim scaling performance on Perlmutter.
Our tests are running on the Perlmutter HPC system. We use the compute node that each contains 4 A100 40GB GPUs. We scaled from 4 nodes (16 GPUs) up to 1024 nodes (4,096 GPUs). For each node count, we selected benchmarks that would neither exceed the GPU memory limits nor fall below 2GB per GPU memory usage, ensuring meaningful problem sizes for each configuration. In our strong scaling evaluation of TANQ-Sim, Figure 7 demonstrates the performance scaling of simulation time across circuit benchmarks. The results show a great scaling of TANQ-Sim with a median parallel efficiency of 67%.
Figure 8. Weak scaling of TANQ-Sim on Perlmutter.
For weak scaling performance, we used per-GPU memory usage as a metric for problem complexity and measured simulation throughput (gates simulated per second) against the number of nodes (GPUs). Figure 8 presents these results. We observed an anomalous high throughput data point for the 2GB per GPU configuration at 256 GPUs, corresponding to the 17-qubit Quantum Volume circuit. This outlier stems from qv_n17’s unusual gate composition, where 80% are 1-qubit gates, which typically require less computational resources than 2-qubit gates. Despite this anomaly, TANQ-Sim demonstrates excellent weak scaling performance, maintaining consistent throughput for fixed per-GPU problem sizes as GPU count increases.